OpenStudy (anonymous):

a balloon filled with helium has a volume of 30.0L at a pressure of 100mmhgand a temperature of 15.0°C. how many moles of helium does the balloon contain?

OpenStudy (anonymous):

First convert 100mmhg(or Torr) to Kpa (Its easier this way) \[100mmHg \times \frac{ 133.3 Pa }{ 1mmHg}\times \frac{ 1Kpa }{ 1000Pa }=13.32Kpa\] Then convert degrees to kelvin (you just have too) \[15^{o}+273.15K =288.15K\] Use the following equation \[pV = nRT\] Ideal gas equation, where 'p' is pressure, 'V' is volume, 'n' is moles, 'R' is a constant (8.314 in Kpa), 'T' is temperature in kelvins. (note: 'R' constant changes with different pressures). \[n=\frac{ pV }{ RT }\]solve for 'n' (moles) \[n=\frac{ 13.32Kpa \times30L }{ 8.314\times288.15 }=1.669molHe\]This is the answer.